a box contains five red balls, six green balls, and nine yellow balls. suppose you select one ball at random from the box and do not replace it. then you randomly select a second ball. find the probability that both balls selected are red.
Answer by Mathtut(3670)

number of ways red can occur 5
total of possible outcomes is 20
:
so to choose 1 red P(r)= 5/20
then you lose one red ball so our number of ways red can occure is now 4 and the total possible outcomes is 19. hence:
:
P(r)=4/19
:
for the probability of this you multiply these together
:
(5/20)(4/19)=(20/20*19)=1/19